OpenStudy (anonymous):

Help me simplify! (-2q^4/7)^4

OpenStudy (anonymous):

x=mc

OpenStudy (anonymous):

I thought x=mc SQUARED?

OpenStudy (anonymous):

x= love y=you

OpenStudy (anonymous):

Aw, how sweet~

OpenStudy (anonymous):

(-2q^4/7)^4 this means that every number inside the parenthesis is being raised to the 4th power, so you would have the following: ((-2)^4*q^8)/(7^4) that gives you 16q^8/2401 because 2401 is odd, and the only multiples of 16 (besides 1) are even (2,4,8), that cannot simplified any further.

OpenStudy (anonymous):

Thank you! You are a god of mathematics.

OpenStudy (anonymous):

Not sure if sarcasm or not...lol. but thanks, i think :)

OpenStudy (anonymous):

No that was seriousness... Because I don't understand hardly anything about math so that was amazing to me. xD

OpenStudy (unklerhaukus):

was you question like this \[\left(\frac{-2q^4}7\right)^4\] or like this \[\left(-2q^{4/7}\right)^4\]?